Please, Please help, As soon as possible!
Use Elimination to solve--
2x=6
x=5-y

Answer by faceoff57(108)

Please, Please help, As soon as possible!
Use Elimination to solve--
2x=6
x=5-y
2(5-y)=6
10-2y=6
10-6=2y
4=2y
2=y
X=5-2
X=3
(3,2)